OpenStudy (cenaida):

Find the slope of the line that contains the following pair of points. (2, 4), (0, 8) A. -2 B. -1/2 C. 1/6 D. 6

OpenStudy (zale101):

Do you remember the slope formula?

OpenStudy (cenaida):

yes slope=vertical change/horizontal change

OpenStudy (zale101):

\(slope=\large \frac{y_2-y_1}{x_2-x_1}\) This is the formula i'm talking about.

OpenStudy (cenaida):

Is it B?

OpenStudy (zale101):

What you do here is take one of the giving point, let's say (2, 4) and plug that for x1,y1 in the formula. So, wherever you see x1, you plug 2, wherever you see y1 you plug 4. Do the same for (0, 8) but it's point will be x2,y2 So, you'll get \(slope=\large\frac{8-4}{0-2}\)

OpenStudy (hockeychick23):

(2, 4), (0, 8) 8-4/0-2=4/-2=?

OpenStudy (cenaida):

-2?

OpenStudy (hockeychick23):

yepp!(:

OpenStudy (cenaida):

thanks girl!

OpenStudy (hockeychick23):

anytime!